OpenStudy (anonymous):

given the domain [-3, -2, -1, 0, 1], determine the range of the given relation. y=x^2-4

OpenStudy (anonymous):

your domain is [-3,1] so , \[-3\le x \le 1 \] you have to give the form of y ( range ) elevating to square \[0 \le x ^{2} \le ( the highest of the exponents) = 9\] \[-4 \le x ^{2} - 4 \le 5\] so you range is [-4,5]